Sara wants to have $600,000 in her savings account when she retires, How much must she put in the account now, if the account pays a fixed interest rate of 10%, to ensure that she has $600,000 in 818 years time?

a. $194,247

b. $107,915

c. $114,639

d. $151,081
Answer #1

Answer: Option b is correct. The correct option is $107915
Future value=Present value*(1+interest rate)^Time period
Taking time period as 18 years,interest rate of 10%, and future value of 600000 we get;
600000=x*(1+10%)^18
=>x=600000/600000=x*(1+10%)^18
x=600000/5.559917313=107915.274 or 107915 (Rounded to the nearest whole number)